OpenStudy (anonymous):

Rita's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $55 a day for food and lodging and $0.45 for each mile traveled. Rita drove 300 miles and was reimbursed $2335. Part A: Create an equation that will determine the number of days x on the trip (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did Rita spend on this trip? (1 points)

OpenStudy (anonymous):

I got the equation. I just don't understand B.

OpenStudy (anonymous):

The expression will be 55+0.45x.

OpenStudy (texaschic101):

x stands for days 55x + 0.45(300) = 2335 now solve for x

OpenStudy (anonymous):

40?

OpenStudy (texaschic101):

I don't know...let me see 55x + 135 = 2335 55x = 2335 - 135 55x = 2200 x = 2200/55 x = 40 yes....40 days

OpenStudy (anonymous):

Oh, okay, thanks. :) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) what does that mean?

OpenStudy (texaschic101):

algebra properties....like distribution property, addition property, multiplication property...etc...

OpenStudy (texaschic101):

do you need help with this part ?

OpenStudy (anonymous):

Yes, I do. :) I don't know how to show that with the expression?

OpenStudy (texaschic101):

55x + 0.45(300) = 2335 -- distributive property of equality 55x + 135 = 2335 -- subtraction property of equality 55x + 135 - 135 = 2335 - 135 -- combine like terms 55x = 2200 -- division property of equality 55x/55 = 2200/55 x = 40

OpenStudy (anonymous):

Oh, thanks sooo much! I get it. :)

OpenStudy (texaschic101):

thats good to hear ....glad to help :)

OpenStudy (anonymous):

Haha...thx again!